OpenStudy (mathmath333):

If u form a subset of integers chosen from between 1 to 3000, such that no two integers add up to a multiple of nine, what can be the maximum numbers of elements in the subset?

OpenStudy (asevilla5):

See, We can write all the numbers from 1 to 3000 as - 9k, (9k+1), (9k+2) and so till (9k+9). Now if we divide 3000 by 9 - Quotient will be 333. Now as the question says - Between 1 and 3000 so exlude them :) Thus numbers having the form - 9k+1 = 333 + 2998 - Number 1 which we have counted and had to be excluded. 9k+2 = 333 + 2999 Similarly, 9k+3, 9k+4, 9k+5, 9k+6, 9k+7, 9k+8, 9k+9 = 333 (3000 Excluded) Now what we have to do here is to choose a subset of integers such that no two integers add up to a multiple of nine. Hence, we cannot take 9k+9. Also, we cannot count in these pairs - (9k+1 and 9k+8 ) (9k+2 and 9k+7 ) (9k+3 and 9k+6 ) (9k+4 and 9k+5 ) as all will be divisible by 9. Now we have to find the maximum number of elements in the set, for that we can calculate it like this - Elements of the form (9k+1 or 9k+ + Elements of the form (9k+2 or 9k+7) + Elements of the form (9k+3 or 9k+6) + Elements of the form (9k+4 or 9k+5) = 333 + 334 + 333 + 333 = 1333 :)

OpenStudy (asevilla5):

so 1333 should be the answer

OpenStudy (rational):

@Asevilla5 's solution makes sense right ?

OpenStudy (anonymous):

so let \(S=\left\{x: 3000>x >1\right\} \) \(u\subseteq S \) with condition :- \( x=x_1\times 10^1+x_2\times 10^2+x_3\times10^3+x_4\times 10^4\) wait lol , i thought u wanna no tow digits hehe nut integers in the subset >.<

OpenStudy (rational):

what are u trying to do @BSwan

OpenStudy (anonymous):

forget , i first missunderstood the question , i thought it like this :- If u form a subset of integers chosen from between 1 to 3000, such that no two digits of the integer add up to a multiple of nine.

OpenStudy (asevilla5):

I tried working out this problem on paper kept messing up and found that question on the website @mathmath333

OpenStudy (rational):

ohkay that looks like a more tough problem

OpenStudy (asevilla5):

mentioned

OpenStudy (anonymous):

not really rashh bhahahaha , however nice try @Asevilla5 ! i like what you did

OpenStudy (mathmath333):